3^{rd} Conference

GEOMETRY AND TOPOLOGY OF MANIFOLDS

Krynica 29 IV – 05 V 2001

http://im0.p.lodz.pl/konferencje/krynica2001

Organized by

– Institute of Mathematics of the Technical University of Lodz Address: Al. Politechniki 11, 90-924 Łódź Jan Kubarski (chairman), e-mail: kubarski@ck-sg.p.lodz.pl

– Institute of Mathematics of the Jagiellonian University, Cracow Address: ul. Wł. Reymonta 4, 30-059 Kraków Robert Wolak, e-mail: wolak@im.uj.edu.pl

– Faculty of Applied Mathematics of the University Mining and Metallurgy, Cracow Address: al. Mickiewicza 30, 30-059 Kraków, Tomasz Rybicki, e-mail: tomasz@uci.agh.edu.pl

THIRD ANNOUNCEMENT

Dear Colleague The day 29 April is destined for the arrival. The University coach will leave the Institute of Mathematics of the Jagiellonian University at 3 p.m. : Wl. Reymonta str. 4, phone: +(48)012-6339781. Please write a short e-mail with the information about your intentions like using the Bus, immediately to Robert Wolak

wolak@im.uj.edu.pl. From the railway station Kraków Główny you should take bus no 179, get off at the fifth stop or tram no 15 and get off at the fifth stop at the corner of Oleandry str. At the end of this street (about 200 m) there is the building of the Institute. The ticket can be bought in kiosks (2.20 PLN and in tram 2.70 PLN). Our first common meal will be the dinner at 7 p.m. at the ”Baśka” hotel in Krynica (phone: +(48)018 4715888) which is situated on a hill overlooking the railway station in Krynica. Kraków can be reached easily from Warszawa by fast trains which leave Warszawa Centralna railway station at 6.15, 6.50, 7.35, 8.35, 11.35, 12.35, 13.35, 15.35, 16.35, 17.35, 19.35 and the last normal train at 21.05 (arrival at 1.54 am). From the Airport to the railway station, you can take a special bus LOT-CENTRUM. The journey, without any stop, takes about 2.40 h, unfortunately the reservation can be obligatory. There are two direct trains from Warszawa to Krynica dep. Warszawa Centralna 6:00, arr. Krynica 12:54; dep. Warszawa Zachodnia 21:11, arr. Krynica 7:02. Sleeping-cars should be also available. It is possible to come to Kraków from the airport by POLSKI EKSPRESS bus, dep. at 9.00 and 15.00 (about 30 PLN). It takes about 7 or 8 hours. In the case of any problems please contact Robert Wolak at wolak@im.uj.edu.pl. Car way through Bochnia and Nowy Sa̧cz. Guests from abroad are requested for pointing the place, day and hour of the arrival to Poland because of individual further way planning:

wolak@im.uj.edu.pl. If you have not sent yet the full registration, the title of your lecture or the abstract or the paper please do it as quickly as possible so we would be able to print materials and present it in the Internet. We plan to publish the proceedings (reviewed and original works) in the Journal ”Universitatis Iagellonicae Acta Mathematica”. Details will be revealed on the Conference.

Best regards and Happy Easters, organizers:

Jan Kubarski, Robert Wolak, Tomasz Rybicki